### v1

# Ultra-low value for the allowed decay of 110Agm confirmed via mass measurements

Abstract
The mass of 109Ag isotope was measured with respect to 110Cd with the phase-imaging ion-cyclotron-resonance technique using the JYFLTRAP double Penning trap at the IGISOL facility. By combining this new result with the known spectroscopic information, the value of eV between the isomer of 110Ag and the state in 110Cd was determined. This is the lowest for any allowed transition observed to date. The nuclear shell-model results were performed with the kshell code employing the jj45pnb interaction, and combined with the state-of-the-art atomic calculations. The theoretical partial half-life, years, and the resulting branching ratio, , combined with a viable production method, thermal-neutron capture on stable 109Ag, make 110Agm as a promising candidate for future antineutrino mass measurement experiments.
The determination of the absolute mass of neutrinos remains one of the big unanswered questions in physics. Currently, the most stringent limits come from the kinematic experiments: KATRIN [1, 2] and MARE [3, 4] for the electron antineutrino, and ECHo [5, 6] and HOLMES [7, 8] for the electron neutrino. They are based on the model-independent analysis of the end-point spectrum of decays of 3H (KATRIN) and 187Re (MARE) or 163Ho electron capture (ECHo and HOLMES). Another tritium-based experiment, Project 8, is currently being planned [9]. These three isotopes have been chosen due to the very small ground-state-to-ground-state values ( eV for 3H [10], eV for 163Ho [11] and eV for 187Re [12]). Nevertheless, despite such narrow energy windows, only a very small fraction of decays ends up near the endpoint, limiting the sensitivity of these experiments.
An alternative approach has been recently proposed where the decay proceeds to the excited state in the daughter isotope with an ultra-low value, usually keV [13]. So far, the most promising candidate is 115In where the second-forbidden unique decay to the first excited states in 115Sn () was experimentally observed [14, 15] and its value is around 200 eV [16, 14]. Two other isotopes of interest are 135Cs where keV for the first-forbidden unique decay to the excited state [17] as well as 131I with the value of keV for an allowed decay to the excited state. [18]. For both 135Cs and 131I, these ultra-low- decays are still awaiting experimental confirmation.
While 115In seems to be an ideal candidate for a new generation of neutrino-mass experiments, a small branching ratio () resulting in a very long partial half-life ( y [14]) is limiting its use. As a result, the search for other cases continues [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30], with an emphasis on allowed transitions as they are the fastest and their shape can be described analytically [31]. However, these studies are focused only on ground-state-to-excited-state decays [31], disregarding possible isomeric-state-to-excited-state (iso-ex) transitions.
One of isomers which might be of interest for the future antineutrino mass experiments is 110Agm ( keV, , d [32, 33]). It decays almost exclusively (98.67(8)%) via to the excited states in stable 110Cd (see Fig. 1) [32, 33] and it can be easily produced in nuclear reactors thanks to the large cross section ( b) for the thermal-neutron capture on stable 109Ag [34, 35]. The literature value is 3008.3(13) keV [36, 32] and the recent studies constrained the excitation energy of the state to keV [37], resulting in the present energy window for the allowed iso-ex decay of keV. Additionally, the state is not isomeric as it deexcites via prompt rays [38, 39]. Consequently, 110Agm is a perfect candidate for a potential ultra-low- transition.
The masses of 109Ag and 110Ag isotopes are connected via a precise single-neutron separation energy value from the study ( keV [40]) while the 110Agm and excitation energies are known from -ray spectroscopy studies [33]. As a result, can be expressed as a function of masses of two stable isotopes, 109Ag and 110Cd, and precise spectroscopic observables. Considering that the quadratically-added uncertainties of , and are equal to 127 eV, the keV uncertainty of [36, 32] can be reduced if the single-proton separation energy () of 110Cd is known with a precision better than 100 eV. In order to accomplish it, the mass of 109Ag was determined with 110Cd as a reference by utilizing the JYFLTRAP double Penning trap mass spectrometer [41] at the Ion Guide Isotope Separator On-Line (IGISOL) facility at the University of Jyväskylä [42]. Additionally, to check for possible mass-dependent systematic effects caused by the u mass difference, two measurements using zirconium oxide molecules, 90Zr16O and 92Zr16O against 91Zr16O, were also conducted.
The silver ions were produced by a discharge ion source placed in the target area and extracted using a sextupole ion guide [43] while the ions of cadmium and zirconium were produced by an offline discharge ion source [44]. After extraction, the ions were accelerated by a 30 kV electrostatic potential and transported through a 55∘ dipole magnet. Then, the continuous beam was sent into the gas-filled radio-frequency quadrupole cooler-buncher [45]. In there, in addition to cooling and bunching the beam, the zirconium oxide molecules were created. Subsequently, the bunches were injected into the first trap of JYFLTRAP, where the axial motion of the bunch was damped by collisions with the gas and mass-selective buffer-gas cooling technique [46] applied to select the ions of interest.
In the second trap, the phase-imaging ion-cyclotron-resonance (PI-ICR) technique [47, 48, 49, 50] was used to measure the ions’ free space cyclotron frequency , where and are the charge and mass of the ion and the magnetic field strength. Two timing patterns were used to excite the ions’ radial eigenmotions. For the magnetron motion, the pattern started with a dipolar radiofrequency pulse exciting the reduced cyclotron motion. This was followed by a quadrupolar pulse converting the motion to magnetron motion after an accumulation time . Finally, the ions were projected onto a position-sensitive microchannel plate detector (2D-MCP). For the reduced cyclotron motion, the order of the converting pulse and the accumulation time was reversed. Typical phase spots obtained in this work are shown in Fig. 2a) and b).
The cyclotron frequency was determined using the angle between the phase spots after magnetron () and cyclotron () motion accumulation, and the number of revolutions the ions performed during :
The mass difference of an ion-of-interest and a reference was determined from the ratio of their measured frequencies.
To reduce the effect of residual magnetron and cyclotron motion on the ions’ position, the start of the conversion pulse was scanned over the cyclotron period and the extraction from the measurement trap over the period of the magnetron motion. The effect of the temporal magnetic field fluctuation was minimized by alternating the measurements of the ion of interest and the reference ion every seven minutes as the fluctuation of the JYFLTRAP magnetic field has been measured to be min-1 [50]. By performing count-rate class analysis [51, 52, 50] using bunches with up to five detected ions, the effect of ion-ion interaction was determined to be negligible.
For the mass determination of 109Ag, a total of 157 frequency ratios between 110Cd and 109Ag were measured over 40 hours, see Fig. 2c). To account for other sources of uncertainties, the data were collected using three different accumulation times, 813 ms, 1 s and 1.5 s, and two different spot positions on the detector. However, no significant effects were observed based on these factors. A final ratio was obtained as a weighted average of all the measurements. Similarly, for 92Zr16O mass determination, 98 ratios over 25 hours were measured using two different spot positions and two accumulation times of 813 ms and 1 s, while for 90Zr16O, 20 ratios with s were collected over five hours, both against the 91Zr16O reference, see Supplemental Material [53].
The single-nucleon separation energy can be determined from the frequency ratio measured between two neighboring species:
where and are the atomic mass of the reference species and the mass of an electron, respectively, while and are the first ionization energies of the reference ion and the ion of interest, respectively, taken from Ref. [54]. For a single-neutron separation energy, is the atomic mass of a neutron and for a single-proton separation energy represents the mass of a hydrogen atom.
| Nuclide | Ref. | (keV) | (keV) | |
| 109Ag | 110Cd | 8918.07(5) | 8917.5(13) | |
| 92Zr16O | 91Zr16O | 8634.74(5) | 8634.75(4) | |
| 90Zr16O | 91Zr16O | 7194.75(12) | 7194.36(15)111From AME20 [36]. | |
| 7194.9(4)222From the ENSDF evaluation [55] based on the study from Ref. [57]. |
The measured frequency ratios and corresponding single-particle separation energies are presented in Table 1. The 92Zr single-neutron separation energy perfectly agrees with the Atomic Mass Evaluation 2020 (AME20) which is based on the study from Ref. [56]. At the same time, the value differs by 0.38(19) keV from AME20. However, it agrees with a value from the measurement [57] evaluated at ENSDF [55]. Considering that between the ion of interest and the reference ion there was only one mass unit of difference, the same as between 109Ag and 110Cd, and that the ZrO molecules are in the same mass region, no additional mass-dependent uncertainties are added to the 109Ag frequency ratio.
The updated keV and the mass excess of 109Ag (ME = keV) agree with AME20 ( keV, ME = keV [36]) and they are 26 and three times more precise, respectively, with the uncertainty of the latter being dominated by the uncertainty of the 110Cd mass. These results allows for reevaluation of the value:
where is the frequency ratio between 110Cd and 109Ag from this work, u is the atomic mass of 110Cd [36], and are the masses of an electron and a neutron, respectively, keV [32] and keV [40] are the 110Agm excitation energy and the 110Ag single-neutron separation energy, respectively, while eV and eV are the first ionization energies of Cd and Ag, respectively [54].
| Transition | AME20 | This work |
| keV | keV | |
| keV | keV | |
| keV | eV |
All the -decay energies extracted in this work agree with AME20 [36] but they are an order of magnitude more precise, see Table 2. The updated is positive, eV, at level. It is the lowest observed value for an allowed transition and the second-lowest after the second-forbidden unique decay of 115In to the first excited state in 115Sn [31].
To estimate the partial half-lives for the decays of 110Ag to the excited states in 110Cd, we utilized the kshell code [58] within the Nuclear Shell Model (NSM) framework, employing the jj45pnb interaction [59], a two-nucleon potential based on a perturbative G-matrix approach. The model space includes the proton orbitals , , , and with single-particle energies of , , , and MeV, respectively. For neutrons, the model space comprises the , , , , and orbitals, with corresponding single-particle energies of , , , , and MeV, as specified in the jj45pnb interaction. The computations were performed without truncation of the model space. We note that for the applied region of , only variations of the jj45pnb interaction are available, yielding similar (or redundant) results within the available Hamiltonians found in codes such as kshell or NuShellX@MSU [60].
| Transition | (keV) | ||||||
| Exp. | 0.5 | 0.6 | 0.7 | 0.8 | |||
| (d) | 808.51 | 561.61 | 390.007 | 286.536 | 219.379 | 0.0117 | 528.88 |
| (d) | 369.02 | 117.627 | 81.6857 | 60.0139 | 45.9482 | -0.4742 | 82.07 |
| (s) | 25.58 | 21.85 | 15.18 | 11.15 | 8.54 | 0.7271 | 2891.22 |
| ( y) | 111This transition has not been observed experimentally yet. | 2.2309 | 1.5492 | 1.1382 | 0.8714 | 0.1542 | 0.405 |
To validate the applied interaction, we computed the excitation energies, magnetic dipole and electric quadrupole moments for both 110Ag and 110Cd, and compared them with evaluated data, see Supplemental Material [53]. The results demonstrated reasonable agreement, except for the quadrupole moment of the 110Ag state. We note that we used standard effective charges: for protons and for neutrons, and the bare g-factors , , , and , and did not attempt to adjust them to reproduce the electromagnetic moments values accurately.
The major known allowed -decay transitions from 110Ag to 110Cd as well as the transition of interest were computed and compared with experimental results, see Table 3. These calculations included screening, radiative, and atomic exchange corrections, originally derived for allowed decays and shown to have a stronger contribution at low electron energies [61]. Additionally, the effective weak axial coupling () was varied between 0.5 and 0.8 as suggested in Ref. [62] for this mass region.
Since the phase-space factor is extremely sensitive to the decay energy, for the transition the expected half-lives for the different values and for the energies ranging from 225 to 600 eV are plotted, see Fig. 3. The experimental uncertainties found in this work are indicated by a gray shaded region. The partial half-life ranges from y for the upper limit (540 eV) with up to y for the lower limit (270 eV) with . For any given curve, we mostly observe an order of magnitude difference in whilst considering the gray shaded area. The partial half-lives for two values, 0.5 and 0.8, were also computed without incorporating the atomic exchange corrections [61] to assess its impact. As can be seen in Fig. 3, the values increase by a factor , and they range between and years. This result highlights the importance of this correction for the ultra-low values.
The theoretical half-lives for the strongest -decay branches are best reproduced when , as shown in Table 3. Therefore, this value was employed for the calculations concerning the main transition of interest. This choice is motivated entirely by its better overall agreement with experimental results, although it reflects a broader overreach of the nuclear shell model, suggesting that the model might not fully capture the quenching effects observed in experimental data. To estimate the influence of the energy window, the experimental value was varied by . The resulting partial half-life is years, which corresponds to a branching ratio of for the decay.
In summary, we have measured with high precision the single-proton separation energy of 110Cd, keV, resulting in the improved mass-excess value of 109Ag, keV. By combining our results with the well known single-neutron separation energy and the isomer excitation energy of 110Ag, we demonstrated that the -decay transition has the ultra-low value of 405(135) eV. This is the lowest for any allowed transition measured to date and the first low case measured for the -decaying isomer. The nuclear shell model combined with the atomic calculations predicted the partial half-life for this transition to be years which translates to the branching ratio of . Consequently, our work indicates that 110Agm might be a good candidate for the future antineutrino mass measurement experiments. It also demonstrates that the future search of low- transitions should not be limited to the -decaying ground states, with the isomers in 101,102Rh, 106Ag, 120Sb, 121Te, 184Re and 194Ir being the most promising cases to find such transitions.
In passing we note that the hints of the decay have been already observed. In two 110Agm decay studies, the 846-keV ray, which is the strongest transition deexciting the state in 110Cd [37], was detected [63, 64]. In addition, a second transition deexciting this level, the 1466-keV ray, was observed in one of these studies [63]. Both of them remain unplaced [63, 64, 33]. However, it should be noted that if these rays originate from the deexcitation of the state, the partial decay half-life to this state would be in the years range, which is orders of magnitude shorter than indicated by calculations. Thus, a future dedicated study to search for the decay is needed.
Acknowledgements.
This project has received funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement No. 771036 (ERC CoG MAIDEN), from the European Union’s Horizon Europe Research and Innovation Programme under Grant Agreement No. 101057511 (EURO-LABS) and the Research Council of Finland projects No. 295207, 306980, 327629, 354589 and 354968. J.R. acknowledges financial support from the Vilho, Yrjö and Kalle Väisälä Foundation.References
- The KATRIN collaboration et al. [2021] The KATRIN collaboration, M. Aker, K. Altenmüller, J.F. Amsbaugh, M. Arenz, M. Babutzka, J. Bast, S. Bauer, H. Bechtler, M. Beck, A. Beglarian, J. Behrens, B. Bender, R. Berendes, A. Berlev, U. Besserer, C. Bettin, B. Bieringer, K. Blaum, F. Block, S. Bobien, M. Böttcher, J. Bohn, K. Bokeloh, H. Bolz, B. Bornschein, L. Bornschein, H. Bouquet, N.M. Boyd, T. Brunst, T.H. Burritt, T.S. Caldwell, Z. Chaoui, S. Chilingaryan, W. Choi, T.J. Corona, G.A. Cox, K. Debowski, M. Deffert, M. Descher, D. Díaz Barrero, P.J. Doe, O. Dragoun, G. Drexlin, J.A. Dunmore, S. Dyba, F. Edzards, F. Eichelhardt, K. Eitel, E. Ellinger, R. Engel, S. Enomoto, M. Erhard, D. Eversheim, M. Fedkevych, A. Felden, S. Fischer, J.A. Formaggio, F.M. Fränkle, G.B. Franklin, H. Frenzel, F. Friedel, A. Fulst, K. Gauda, R. Gehring, W. Gil, F. Glück, S. Görhardt, J. Grimm, R. Grössle, S. Groh, S. Grohmann, R. Gumbsheimer, M. Hackenjos, D. Häßler, V. Hannen, F. Harms, G.C. Harper, J. Hartmann, N. Haußmann, F. Heizmann, K. Helbing, M. Held, S. Hickford, D. Hilk, B. Hillen, R. Hiller, D. Hillesheimer, D. Hinz, T. Höhn, M. Hötzel, S. Holzmann, S. Horn, T. Houdy, M.A. Howe, A. Huber, T. James, A. Jansen, M. Kaiser, C. Karl, O. Kazachenko, J. Kellerer, L. Kippenbrock, M. Kleesiek, M. Kleifges, J. Kleinfeller, M. Klein, C. Köhler, L. Köllenberger, A. Kopmann, M. Korzeczek, A. Kosmider, A. Kovalík, B. Krasch, H. Krause, M. Kraus, L. Kuckert, A. Kumb, N. Kunka, T. Lasserre, L. La Cascio, O. Lebeda, M.L. Leber, B. Lehnert, B. Leiber, J. Letnev, R.J. Lewis, T.L. Le, S. Lichter, A. Lokhov, J.M. Lopez Poyato, M. Machatschek, E. Malcherek, M. Mark, A. Marsteller, E.L. Martin, K. Mehret, M. Meloni, C. Melzer, A. Menshikov, S. Mertens, L.I. Minter (née Bodine), B. Monreal, J. Mostafa, K. Müller, A.W. Myers, U. Naumann, H. Neumann, S. Niemes, P. Oelpmann, A. Off, H.-W. Ortjohann, A. Osipowicz, B. Ostrick, D.S. Parno, D.A. Peterson, P. Plischke, A.W.P. Poon, M. Prall, F. Priester, P.C.-O. Ranitzsch, J. Reich, P. Renschler, O. Rest, R. Rinderspacher, R.G.H. Robertson, W. Rodejohann, C. Rodenbeck, M. Röllig, C. Röttele, P. Rohr, S. Rupp, M. Ryšavý, R. Sack, A. Saenz, M. Sagawe, P. Schäfer, A. Schaller (née Pollithy), L. Schimpf, K. Schlösser, M. Schlösser, L. Schlüter, S. Schneidewind, H. Schön, K. Schönung, M. Schrank, B. Schulz, J. Schwarz, M. Šefčík, H. Seitz-Moskaliuk, W. Seller, V. Sibille, D. Siegmann, M. Slezák, F. Spanier, M. Steidl, M. Sturm, M. Sun, D. Tcherniakhovski, H.H. Telle, L.A. Thorne, T. Thümmler, N. Titov, I. Tkachev, N. Trost, K. Urban, K. Valerius, B.A. VanDevender, T.D. Van Wechel, D. Vénos, A. Verbeek, R. Vianden, A.P. Vizcaya Hernández, K. Vogt, B.L. Wall, N. Wandkowsky, M. Weber, H. Weingardt, C. Weinheimer, C. Weiss, S. Welte, J. Wendel, K.J. Wierman, J.F. Wilkerson, J. Wolf, S. Wüstling, W. Xu, Y.-R. Yen, M. Zacher, S. Zadoroghny, M. Zboril, and G. Zeller, The design, construction, and commissioning of the KATRIN experiment, Journal of Instrumentation 16 (08), T08015.
- The KATRIN Collaboration et al. [2022] The KATRIN Collaboration, M. Aker, A. Beglarian, J. Behrens, A. Berlev, U. Besserer, B. Bieringer, F. Block, S. Bobien, M. Böttcher, B. Bornschein, L. Bornschein, T. Brunst, T. S. Caldwell, R. M. D. Carney, L. La Cascio, S. Chilingaryan, W. Choi, K. Debowski, M. Deffert, M. Descher, D. Díaz Barrero, P. J. Doe, O. Dragoun, G. Drexlin, K. Eitel, E. Ellinger, R. Engel, S. Enomoto, A. Felden, J. A. Formaggio, F. M. Fränkle, G. B. Franklin, F. Friedel, A. Fulst, K. Gauda, W. Gil, F. Glück, R. Grössle, R. Gumbsheimer, V. Gupta, T. Höhn, V. Hannen, N. Haußmann, K. Helbing, S. Hickford, R. Hiller, D. Hillesheimer, D. Hinz, T. Houdy, A. Huber, A. Jansen, C. Karl, F. Kellerer, J. Kellerer, M. Kleifges, M. Klein, C. Köhler, L. Köllenberger, A. Kopmann, M. Korzeczek, A. Kovalík, B. Krasch, H. Krause, N. Kunka, T. Lasserre, T. L. Le, O. Lebeda, B. Lehnert, A. Lokhov, M. Machatschek, E. Malcherek, M. Mark, A. Marsteller, E. L. Martin, C. Melzer, A. Menshikov, S. Mertens, J. Mostafa, K. Müller, H. Neumann, S. Niemes, P. Oelpmann, D. S. Parno, A. W. P. Poon, J. M. L. Poyato, F. Priester, S. Ramachandran, R. G. H. Robertson, W. Rodejohann, M. Röllig, C. Röttele, C. Rodenbeck, M. Ryšavý, R. Sack, A. Saenz, P. Schäfer, A. Schaller née Pollithy, L. Schimpf, K. Schlösser, M. Schlösser, L. Schlüter, S. Schneidewind, M. Schrank, B. Schulz, A. Schwemmer, M. Šefčík, V. Sibille, D. Siegmann, M. Slezák, F. Spanier, M. Steidl, M. Sturm, M. Sun, D. Tcherniakhovski, H. H. Telle, L. A. Thorne, T. Thümmler, N. Titov, I. Tkachev, K. Urban, K. Valerius, D. Vénos, A. P. Vizcaya Hernández, C. Weinheimer, S. Welte, J. Wendel, J. F. Wilkerson, J. Wolf, S. Wüstling, J. Wydra, W. Xu, Y.-R. Yen, S. Zadoroghny, and G. Zeller, Direct neutrino-mass measurement with sub-electronvolt sensitivity, Nature Physics 18, 160 (2022).
- Nucciotti [2008] A. Nucciotti, The MARE Project, Journal of Low Temperature Physics 151, 597 (2008).
- Ferri et al. [2015] E. Ferri, D. Bagliani, M. Biasotti, G. Ceruti, D. Corsini, M. Faverzani, F. Gatti, A. Giachero, C. Gotti, C. Kilbourne, A. Kling, M. Maino, P. Manfrinetti, A. Nucciotti, G. Pessina, G. Pizzigoni, M. Ribeiro Gomes, and M. Sisti, The Status of the MARE Experiment with 187Re and 163Ho Isotopes, Physics Procedia 61, 227 (2015).
- Gastaldo et al. [2017] L. Gastaldo, K. Blaum, K. Chrysalidis, T. Day Goodacre, A. Domula, M. Door, H. Dorrer, C. E. Düllmann, K. Eberhardt, S. Eliseev, C. Enss, A. Faessler, P. Filianin, A. Fleischmann, D. Fonnesu, L. Gamer, R. Haas, C. Hassel, D. Hengstler, J. Jochum, K. Johnston, U. Kebschull, S. Kempf, T. Kieck, U. Köster, S. Lahiri, M. Maiti, F. Mantegazzini, B. Marsh, P. Neroutsos, Y. N. Novikov, P. C. O. Ranitzsch, S. Rothe, A. Rischka, A. Saenz, O. Sander, F. Schneider, S. Scholl, R. X. Schüssler, C. Schweiger, F. Simkovic, T. Stora, Z. Szücs, A. Türler, M. Veinhard, M. Weber, M. Wegner, K. Wendt, and K. Zuber, The electron capture in 163Ho experiment – ECHo, The European Physical Journal Special Topics 226, 1623 (2017).
- Velte et al. [2019] C. Velte, F. Ahrens, A. Barth, K. Blaum, M. Braß, M. Door, H. Dorrer, C. E. Düllmann, S. Eliseev, C. Enss, P. Filianin, A. Fleischmann, L. Gastaldo, A. Goeggelmann, T. D. Goodacre, M. W. Haverkort, D. Hengstler, J. Jochum, K. Johnston, M. Keller, S. Kempf, T. Kieck, C. M. König, U. Köster, K. Kromer, F. Mantegazzini, B. Marsh, Y. N. Novikov, F. Piquemal, C. Riccio, D. Richter, A. Rischka, S. Rothe, R. X. Schüssler, C. Schweiger, T. Stora, M. Wegner, K. Wendt, M. Zampaolo, and K. Zuber, High-resolution and low-background 163Ho spectrum: interpretation of the resonance tails, The European Physical Journal C 79, 1026 (2019).
- Faverzani et al. [2016] M. Faverzani, B. Alpert, D. Backer, D. Bennet, M. Biasotti, C. Brofferio, V. Ceriale, G. Ceruti, D. Corsini, P. K. Day, M. De Gerone, R. Dressler, E. Ferri, J. Fowler, E. Fumagalli, J. Gard, F. Gatti, A. Giachero, J. Hays-Wehle, S. Heinitz, G. Hilton, U. Köster, M. Lusignoli, M. Maino, J. Mates, S. Nisi, R. Nizzolo, A. Nucciotti, A. Orlando, L. Parodi, G. Pessina, G. Pizzigoni, A. Puiu, S. Ragazzi, C. Reintsema, M. Ribeiro-Gomez, D. Schmidt, D. Schuman, F. Siccardi, M. Sisti, D. Swetz, F. Terranova, J. Ullom, and L. Vale, The HOLMES Experiment, Journal of Low Temperature Physics 184, 922 (2016).
- De Gerone et al. [2022] M. De Gerone, B. Alpert, M. Balata, D. T. Becker, D. A. Bennett, A. Bevilacqua, M. Borghesi, G. Ceruti, G. D. B. De Galembert, R. Dressler, M. Faverzani, M. Fedkevych, E. Ferri, J. W. Fowler, G. Gallucci, J. D. Gard, F. Gatti, A. Giachero, G. C. Hilton, U. Köster, M. Lusignoli, P. Manfrinetti, J. A. B. Mates, E. Maugeri, S. Nisi, A. Nucciotti, L. Parodi, G. Pessina, S. Ragazzi, C. D. Reintsema, D. R. Schmidt, D. Schumann, F. Siccardi, D. S. Swetz, J. N. Ullom, and L. R. Vale, Status of the HOLMES Experiment, Journal of Low Temperature Physics 209, 980 (2022).
- Pettus and on behalf of theProject 8 Collaboration [2020] W. C. Pettus and on behalf of theProject 8 Collaboration, Overview of Project 8 and Progress Towards Tritium Operation, Journal of Physics: Conference Series 1342, 012040 (2020).
- Medina Restrepo and Myers [2023] M. Medina Restrepo and E. G. Myers, Mass Difference of Tritium and Helium-3, Phys. Rev. Lett. 131, 243002 (2023).
- Schweiger et al. [2024] C. Schweiger, M. Braß, V. Debierre, M. Door, H. Dorrer, C. E. Düllmann, C. Enss, P. Filianin, L. Gastaldo, Z. Harman, M. W. Haverkort, J. Herkenhoff, P. Indelicato, C. H. Keitel, K. Kromer, D. Lange, Y. N. Novikov, D. Renisch, A. Rischka, R. X. Schüssler, S. Eliseev, and K. Blaum, Penning-trap measurement of the value of electron capture in 163Ho for the determination of the electron neutrino mass, Nature Physics 20, 921 (2024).
- Filianin et al. [2021] P. Filianin, C. Lyu, M. Door, K. Blaum, W. J. Huang, M. Haverkort, P. Indelicato, C. H. Keitel, K. Kromer, D. Lange, Y. N. Novikov, A. Rischka, R. X. Schüssler, C. Schweiger, S. Sturm, S. Ulmer, Z. Harman, and S. Eliseev, Direct -Value Determination of the Decay of , Phys. Rev. Lett. 127, 072502 (2021).
- Kopp and Merle [2010] J. Kopp and A. Merle, Ultralow values for neutrino mass measurements, Phys. Rev. C 81, 045501 (2010).
- Wieslander et al. [2009] J. S. E. Wieslander, J. Suhonen, T. Eronen, M. Hult, V.-V. Elomaa, A. Jokinen, G. Marissens, M. Misiaszek, M. T. Mustonen, S. Rahaman, C. Weber, and J. Äystö, Smallest Known Value of Any Nuclear Decay: The Rare Decay of , Phys. Rev. Lett. 103, 122501 (2009).
- Andreotti et al. [2011] E. Andreotti, M. Hult, R. González de Orduña, G. Marissens, J. S. E. Wieslander, and M. Misiaszek, Half-life of the decay 115In(9/2+)115Sn(3/2+), Phys. Rev. C 84, 044605 (2011).
- Mount et al. [2009] B. J. Mount, M. Redshaw, and E. G. Myers, Value of : The Lowest Known Energy Decay, Phys. Rev. Lett. 103, 122502 (2009).
- de Roubin et al. [2020] A. de Roubin, J. Kostensalo, T. Eronen, L. Canete, R. P. de Groote, A. Jokinen, A. Kankainen, D. A. Nesterenko, I. D. Moore, S. Rinta-Antila, J. Suhonen, and M. Vilén, High-Precision -Value Measurement Confirms the Potential of for Absolute Antineutrino Mass Scale Determination, Phys. Rev. Lett. 124, 222503 (2020).
- Eronen et al. [2022] T. Eronen, Z. Ge, A. de Roubin, M. Ramalho, J. Kostensalo, J. Kotila, O. Beliushkina, C. Delafosse, S. Geldhof, W. Gins, M. Hukkanen, A. Jokinen, A. Kankainen, I.D. Moore, D.A. Nesterenko, M. Stryjczyk, and J. Suhonen, High-precision measurement of a low Q value for allowed -decay of 131I related to neutrino mass determination, Physics Letters B 830, 137135 (2022).
- Sandler et al. [2019] R. Sandler, G. Bollen, N. D. Gamage, A. Hamaker, C. Izzo, D. Puentes, M. Redshaw, R. Ringle, and I. Yandow, Investigation of the potential ultralow -value -decay candidates and using Penning trap mass spectrometry, Phys. Rev. C 100, 024309 (2019).
- Karthein et al. [2019] J. Karthein, D. Atanasov, K. Blaum, S. Eliseev, P. Filianin, D. Lunney, V. Manea, M. Mougeot, D. Neidherr, Y. Novikov, L. Schweikhard, A. Welker, F. Wienholtz, and K. Zuber, Direct decay-energy measurement as a route to the neutrino mass, Hyperfine Interactions 240, 61 (2019).
- Ge et al. [2021a] Z. Ge, T. Eronen, A. de Roubin, D. A. Nesterenko, M. Hukkanen, O. Beliuskina, R. de Groote, S. Geldhof, W. Gins, A. Kankainen, A. Koszorús, J. Kotila, J. Kostensalo, I. D. Moore, A. Raggio, S. Rinta-Antila, J. Suhonen, V. Virtanen, A. P. Weaver, A. Zadvornaya, and A. Jokinen, Direct measurement of the mass difference of rules out as a promising -decay candidate to determine the neutrino mass, Phys. Rev. C 103, 065502 (2021a).
- Ge et al. [2021b] Z. Ge, T. Eronen, K. S. Tyrin, J. Kotila, J. Kostensalo, D. A. Nesterenko, O. Beliuskina, R. de Groote, A. de Roubin, S. Geldhof, W. Gins, M. Hukkanen, A. Jokinen, A. Kankainen, A. Koszorús, M. I. Krivoruchenko, S. Kujanpää, I. D. Moore, A. Raggio, S. Rinta-Antila, J. Suhonen, V. Virtanen, A. P. Weaver, and A. Zadvornaya, Electron-Capture: A New Candidate for Neutrino Mass Determination, Phys. Rev. Lett. 127, 272301 (2021b).
- Ramalho et al. [2022] M. Ramalho, Z. Ge, T. Eronen, D. A. Nesterenko, J. Jaatinen, A. Jokinen, A. Kankainen, J. Kostensalo, J. Kotila, M. I. Krivoruchenko, J. Suhonen, K. S. Tyrin, and V. Virtanen, Observation of an ultralow--value electron-capture channel decaying to via a high-precision mass measurement, Phys. Rev. C 106, 015501 (2022).
- Ge et al. [2022a] Z. Ge, T. Eronen, A. de Roubin, K.S. Tyrin, L. Canete, S. Geldhof, A. Jokinen, A. Kankainen, J. Kostensalo, J. Kotila, M.I. Krivoruchenko, I.D. Moore, D.A. Nesterenko, J. Suhonen, and M. Vilén, High-precision electron-capture value measurement of 111In for electron-neutrino mass determination, Physics Letters B 832, 137226 (2022a).
- Ge et al. [2022b] Z. Ge, T. Eronen, A. de Roubin, J. Kostensalo, J. Suhonen, D. A. Nesterenko, O. Beliuskina, R. de Groote, C. Delafosse, S. Geldhof, W. Gins, M. Hukkanen, A. Jokinen, A. Kankainen, J. Kotila, A. Koszorús, I. D. Moore, A. Raggio, S. Rinta-Antila, V. Virtanen, A. P. Weaver, and A. Zadvornaya, Direct determination of the atomic mass difference of the pairs and rules out and as possible candidates for electron (anti)neutrino mass measurements, Phys. Rev. C 106, 015502 (2022b).
- Gamage et al. [2022] N. D. Gamage, R. Sandler, F. Buchinger, J. A. Clark, D. Ray, R. Orford, W. S. Porter, M. Redshaw, G. Savard, K. S. Sharma, and A. A. Valverde, Precise -value measurements of and with the Canadian Penning trap for evaluation of potential ultralow -value decays, Phys. Rev. C 106, 045503 (2022).
- Horana Gamage et al. [2022] M. Horana Gamage, R. Bhandari, G. Bollen, N. D. Gamage, A. Hamaker, D. Puentes, M. Redshaw, R. Ringle, S. Schwarz, C. S. Sumithrarachchi, and I. Yandow, Identification of a potential ultralow--value electron-capture decay branch in via a precise Penning trap measurement of the mass of , Phys. Rev. C 106, 065503 (2022).
- Ge et al. [2023] Z. Ge, T. Eronen, A. de Roubin, M. Ramalho, J. Kostensalo, J. Kotila, J. Suhonen, D. A. Nesterenko, A. Kankainen, P. Ascher, O. Beliuskina, M. Flayol, M. Gerbaux, S. Grévy, M. Hukkanen, A. Husson, A. Jaries, A. Jokinen, I. D. Moore, P. Pirinen, J. Romero, M. Stryjczyk, V. Virtanen, and A. Zadvornaya, decay -value measurement of and its implications for neutrino studies, Phys. Rev. C 108, 045502 (2023).
- Ge et al. [2024a] Z. Ge, T. Eronen, M. Ramalho, A. de Roubin, D. A. Nesterenko, A. Kankainen, O. Beliuskina, R. de Groote, S. Geldhof, W. Gins, M. Hukkanen, A. Jokinen, Á. Koszorús, J. Kotila, J. Kostensalo, I. D. Moore, P. Pirinen, A. Raggio, S. Rinta-Antila, V. A. Sevestrean, J. Suhonen, V. Virtanen, and A. Zadvornaya, Direct high-precision measurement of the mass difference of 77As–77Se related to neutrino mass determination, The European Physical Journal A 60, 104 (2024a).
- Ge et al. [2024b] Z. Ge, T. Eronen, V. A. Sevestrean, O. Niţescu, S. Stoica, M. Ramalho, J. Suhonen, A. de Roubin, D. Nesterenko, A. Kankainen, P. Ascher, S. A. S. Andres, O. Beliuskina, P. Delahaye, M. Flayol, M. Gerbaux, S. Grévy, M. Hukkanen, A. Jaries, A. Jokinen, A. Husson, D. Kahl, J. Kostensalo, J. Kotila, I. Moore, S. Nikas, M. Stryjczyk, and V. Virtanen, High-precision measurements of the atomic mass and electron-capture decay value of 95Tc (2024b), arXiv:2406.05273 [nucl-ex] .
- Keblbeck et al. [2023] D. K. Keblbeck, R. Bhandari, N. D. Gamage, M. Horana Gamage, K. G. Leach, X. Mougeot, and M. Redshaw, Updated evaluation of potential ultralow -value -decay candidates, Phys. Rev. C 107, 015504 (2023).
- F.G. Kondev et al. [2021] F.G. Kondev, M. Wang, W.J. Huang, S. Naimi, and G. Audi, The NUBASE2020 evaluation of nuclear physics properties, Chinese Physics C 45, 030001 (2021).
- Gürdal and F.G. Kondev [2012] G. Gürdal and F.G. Kondev, Nuclear Data Sheets for A = 110, Nuclear Data Sheets 113, 1315 (2012).
- Nakamura et al. [2003] S. Nakamura, H. Wada, O. Shcherbakov, K. Furutaka, H. Harada, and T. Katoh, Measurement of the Thermal Neutron Capture Cross Section and the Resonance Integral of the 109Ag(n, )110mAg Reaction, Journal of Nuclear Science and Technology 40, 119 (2003).
- Elmaghraby et al. [2018] E. K. Elmaghraby, E. Salem, Z. Yousef, and N. El-Anwar, Role of isomeric state formation on the measurement of thermal neutron cross section and resonance integral, Physica Scripta 94, 015301 (2018).
- Meng Wang et al. [2021] Meng Wang, W.J. Huang, F.G. Kondev, G. Audi, and S. Naimi, The AME 2020 atomic mass evaluation (II). Tables, graphs and references, Chinese Physics C 45, 030003 (2021).
- [37] B. Singh, Compiled (unevaluated) dataset from Refs. [38] and [39], compiled on 15 Jun 2020, from XUNDL database as of 01 Sep 2024. Version available at https://www.nndc.bnl.gov/xundlarchivals/.
- Garrett et al. [2019] P. E. Garrett, T. R. Rodríguez, A. D. Varela, K. L. Green, J. Bangay, A. Finlay, R. A. E. Austin, G. C. Ball, D. S. Bandyopadhyay, V. Bildstein, S. Colosimo, D. S. Cross, G. A. Demand, P. Finlay, A. B. Garnsworthy, G. F. Grinyer, G. Hackman, B. Jigmeddorj, J. Jolie, W. D. Kulp, K. G. Leach, A. C. Morton, J. N. Orce, C. J. Pearson, A. A. Phillips, A. J. Radich, E. T. Rand, M. A. Schumaker, C. E. Svensson, C. Sumithrarachchi, S. Triambak, N. Warr, J. Wong, J. L. Wood, and S. W. Yates, Multiple Shape Coexistence in , Phys. Rev. Lett. 123, 142502 (2019).
- Garrett et al. [2020] P. E. Garrett, T. R. Rodríguez, A. Diaz Varela, K. L. Green, J. Bangay, A. Finlay, R. A. E. Austin, G. C. Ball, D. S. Bandyopadhyay, V. Bildstein, S. Colosimo, D. S. Cross, G. A. Demand, P. Finlay, A. B. Garnsworthy, G. F. Grinyer, G. Hackman, B. Jigmeddorj, J. Jolie, W. D. Kulp, K. G. Leach, A. C. Morton, J. N. Orce, C. J. Pearson, A. A. Phillips, A. J. Radich, E. T. Rand, M. A. Schumaker, C. E. Svensson, C. Sumithrarachchi, S. Triambak, N. Warr, J. Wong, J. L. Wood, and S. W. Yates, Shape coexistence and multiparticle-multihole structures in , Phys. Rev. C 101, 044302 (2020).
- Bogdanovic et al. [1981] M. Bogdanovic, T.D. Mac Mahon, and H. Seyfarth, Low-Lying States in Doubly-Odd 108Ag and 110Ag, in Annual Report / Kernforschungsanlage Jülich Institut für Kernphysik : 1980, JUL-Spez-99,P76 (Kernforschungsanlage Jülich GmbH Zentralbibliothek, Verlag, 1981) p. 76.
- Eronen et al. [2012] T. Eronen, V. S. Kolhinen, V. V. Elomaa, D. Gorelov, U. Hager, J. Hakala, A. Jokinen, A. Kankainen, P. Karvonen, S. Kopecky, I. D. Moore, H. Penttilä, S. Rahaman, S. Rinta-Antila, J. Rissanen, A. Saastamoinen, J. Szerypo, C. Weber, and J. Äystö, JYFLTRAP: a Penning trap for precision mass spectroscopy and isobaric purification, The European Physical Journal A 48, 46 (2012).
- I.D. Moore et al. [2013] I.D. Moore, T. Eronen, D. Gorelov, J. Hakala, A. Jokinen, A. Kankainen, V.S. Kolhinen, J. Koponen, H. Penttilä, I. Pohjalainen, M. Reponen, J. Rissanen, A. Saastamoinen, S. Rinta-Antila, V. Sonnenschein, and J. Äystö, Towards commissioning the new IGISOL-4 facility, Nuclear Instruments and Methods in Physics Research Section B: Beam Interactions with Materials and Atoms 317, 208 (2013).
- Karvonen et al. [2008] P. Karvonen, I.D. Moore, T. Sonoda, T. Kessler, H. Penttilä, K. Peräjärvi, P. Ronkanen, and J. Äystö, A sextupole ion beam guide to improve the efficiency and beam quality at IGISOL, Nuclear Instruments and Methods in Physics Research Section B: Beam Interactions with Materials and Atoms 266, 4794 (2008).
- Vilén et al. [2020] M. Vilén, L. Canete, B. Cheal, A. Giatzoglou, R. de Groote, A. de Roubin, T. Eronen, S. Geldhof, A. Jokinen, A. Kankainen, I.D. Moore, D.A. Nesterenko, H. Penttilä, I. Pohjalainen, M. Reponen, and S. Rinta-Antila, A new off-line ion source facility at IGISOL, Nuclear Instruments and Methods in Physics Research Section B: Beam Interactions with Materials and Atoms 463, 382 (2020).
- Nieminen et al. [2001] A. Nieminen, J. Huikari, A. Jokinen, J. Äystö, P. Campbell, and E.C.A. Cochrane, Beam cooler for low-energy radioactive ions, Nuclear Instruments and Methods in Physics Research Section A: Accelerators, Spectrometers, Detectors and Associated Equipment 469, 244 (2001).
- Savard et al. [1991] G. Savard, St. Becker, G. Bollen, H.-J. Kluge, R. Moore, T. Otto, L. Schweikhard, H. Stolzenberg, and U. Wiess, A new cooling technique for heavy ions in a Penning trap, Physics Letters A 158, 247 (1991).
- Eliseev et al. [2013] S. Eliseev, K. Blaum, M. Block, C. Droese, M. Goncharov, E. Minaya Ramirez, D. A. Nesterenko, Y. N. Novikov, and L. Schweikhard, Phase-Imaging Ion-Cyclotron-Resonance Measurements for Short-Lived Nuclides, Phys. Rev. Lett. 110, 082501 (2013).
- Eliseev et al. [2014] S. Eliseev, K. Blaum, M. Block, A. Dörr, C. Droese, T. Eronen, M. Goncharov, M. Höcker, J. Ketter, E. M. Ramirez, D. A. Nesterenko, Y. N. Novikov, and L. Schweikhard, A phase-imaging technique for cyclotron-frequency measurements, Applied Physics B 114, 107 (2014).
- Nesterenko et al. [2018] D. A. Nesterenko, T. Eronen, A. Kankainen, L. Canete, A. Jokinen, I. D. Moore, H. Penttilä, S. Rinta-Antila, A. de Roubin, and M. Vilen, Phase-Imaging Ion-Cyclotron-Resonance technique at the JYFLTRAP double Penning trap mass spectrometer, The European Physical Journal A 54, 154 (2018).
- Nesterenko et al. [2021] D. A. Nesterenko, T. Eronen, Z. Ge, A. Kankainen, and M. Vilen, Study of radial motion phase advance during motion excitations in a Penning trap and accuracy of JYFLTRAP mass spectrometer, The European Physical Journal A 57, 302 (2021).
- Kellerbauer et al. [2003] A. Kellerbauer, K. Blaum, G. Bollen, F. Herfurth, H.-J. Kluge, M. Kuckein, E. Sauvan, C. Scheidenberger, and L. Schweikhard, From direct to absolute mass measurements: A study of the accuracy of ISOLTRAP, Eur. Phys. J. D 22, 53 (2003).
- Roux et al. [2013] C. Roux, K. Blaum, M. Block, C. Droese, S. Eliseev, M. Goncharov, F. Herfurth, E. M. Ramirez, D. A. Nesterenko, Y. N. Novikov, and L. Schweikhard, Data analysis of Q-value measurements for double-electron capture with SHIPTRAP, The European Physical Journal D 67, 146 (2013).
- [53] See Supplemental Material at [URL will be inserted by publisher] for an additional Figure with the measured frequency ratios and Tables with the results of Nuclear Shell Model calculations compared with experimental data from Refs. [33, 65, 66, 67, 68, 37].
- Kramida et al. [2022] A. Kramida, Y. Ralchenko, J. Reader, and NIST ASD Team, NIST Standard Reference Database 78, Version 5.10 (2022).
- Baglin [2013] C. M. Baglin, Nuclear Data Sheets for A = 91, Nuclear Data Sheets 114, 1293 (2013).
- Rząca-Urban et al. [2018] T. Rząca-Urban, W. Urban, M. Czerwiński, J. Wiśniewski, A. Blanc, H. Faust, M. Jentschel, P. Mutti, U. Köster, T. Soldner, G. de France, G. S. Simpson, and C. A. Ur, Low-spin excitations in , Phys. Rev. C 98, 064315 (2018).
- Lone and Bartholomew [1979] M. A. Lone and G. A. Bartholomew, 90Zr(n,) Studies with Thermal Neutrons, in Neutron Capture Gamma-Ray Spectroscopy, edited by R. E. Chrien and W. R. Kane (Springer US, Boston, MA, 1979) pp. 675–677.
- Shimizu et al. [2019] N. Shimizu, T. Mizusaki, Y. Utsuno, and Y. Tsunoda, Thick-restart block Lanczos method for large-scale shell-model calculations, Computer Physics Communications 244, 372 (2019).
- Lisetskiy et al. [2004] A. F. Lisetskiy, B. A. Brown, M. Horoi, and H. Grawe, New effective interactions for the model space: Implications for valence-mirror symmetry and seniority isomers, Phys. Rev. C 70, 044314 (2004).
- Brown and Rae [2014] B. A. Brown and W. D. M. Rae, The Shell-Model Code NuShellX@MSU, Nuclear Data Sheets 120, 115 (2014).
- Niţescu et al. [2023] O. Niţescu, S. Stoica, and F. Šimkovic, Exchange correction for allowed decay, Phys. Rev. C 107, 025501 (2023).
- Ejiri et al. [2019] H. Ejiri, J. Suhonen, and K. Zuber, Neutrino–nuclear responses for astro-neutrinos, single beta decays and double beta decays, Physics Reports 797, 1 (2019).
- Mallet [1981] G. Mallet, Measurement of the Radioactive Decay of 110Agm+g, Journal of the Physical Society of Japan 50, 384 (1981).
- Kiang et al. [1993] L. Kiang, P. Teng, G. Kiang, W. Chang, and P. Tu, Study on the Level Structure of 110Cd Populated by the Beta Decay of 110mAg, Journal of the Physical Society of Japan 62, 888 (1993).
- N.J. Stone [2016] N.J. Stone, Table of nuclear electric quadrupole moments, Atomic Data and Nuclear Data Tables 111-112, 1 (2016).
- N.J. Stone [2019] N.J. Stone, Table of Recommended Nuclear Magnetic Dipole Moments: Part I, Long-lived States, Tech. Rep. INDC(NDS)-0794 (International Atomic Energy Agency, 2019).
- N.J. Stone [2020] N.J. Stone, Table of Recommended Nuclear Magnetic Dipole Moments: Part II, Short-lived States, Tech. Rep. INDC(NDS)-0816 (International Atomic Energy Agency, 2020).
- Zhang et al. [2023] H. C. Zhang, K. Y. Ma, J. X. Teng, Z. H. Zhao, H. Wang, S. Y. Liu, Y. C. Hao, J. B. Lu, Y. J. Ma, D. Yang, X. G. Wu, Y. Zheng, and C. B. Li, New positive-parity bands in and systematic studies in silver isotopes, Phys. Rev. C 108, 024314 (2023).